the form of an extra dividend to shareholders or embark on a share repurchase scheme. Current profits are R3.40 per share and their shares currently trade for R35. Their abbreviated balance sheet before paying out the dividend is as follows:

Equity 350 000 Bank/cash 130 000 Debt 120 000 Other assets 340 000 470 000 470 000 Evaluate each alternative(i.e. pay the dividend or repurchase the shares) by:Calculating the price earnings ratio.
1
Expert's answer
2019-09-26T14:35:34-0400

Price earnings ratio P/E ratio = Share price/Earnings per share = 35/3.4 = 10.29.